An investment of Php 100 000 earned a net income of Php 3 450 in one year. Part of the investment was in bonds and earned an income at a rate of 8%. The rest of the investment was in stocks and lost money at a rate of 5%. How much was invested in bonds?

Found 2 solutions by ikleyn, solve_for_x:
Answer by ikleyn(52775)
Answer by solve_for_x(190)

Let B represent the amount invested in bonds, and let S represent the amount invested in stocks.

The sum of the two must equal 100,000, which gives:

B + S = 100,000

The amount of income earned is then:

0.08B - 0.05S = 3450

Solving the first equation for S gives:

S = 100,000 - B

Substituting this into the second equation gives:

0.08B - 0.05(100,000 - B) = 3450

0.08B + 0.05B - 5000 = 3450

0.13B = 3450 + 5000

0.13B = 8450

B = 8450 / 0.13

B = 65,000

The amount invested in bonds was Php 65,000.